# LayoutVLM: Differentiable Optimization of 3D Layout via Vision-Language Models

### v1

@@ Abstract
Open-universe 3D layout generation arranges unlabeled 3D assets conditioned on language instruction. Large language models (LLMs) struggle with generating physically plausible 3D scenes and adherence to input instructions, particularly in dense scenes. We introduce LayoutVLM, a framework and scene layout representation that exploits the semantic knowledge of Vision-Language Models (VLMs) and supports differentiable optimization to ensure physical plausibility. LayoutVLM employs VLMs to generate two mutually reinforcing representations from visually marked images, and a self-consistent decoding process to improve VLMs spatial planning. Our experiments show that LayoutVLM addresses the limitations of existing LLM and constraint-based approaches, producing physically plausible 3D layouts better aligned with the semantic intent of input language instructions. We also demonstrate that fine-tuning VLMs with the proposed scene layout representation extracted from existing scene datasets can improve performance.
@@ 4.1 Scene Layout Representation
Differentiable Spatial Relations. The goal of these spatial relations is twofold: (a) to capture the semantics of the input language instruction and (b) to preserve these semantics during optimization for physical plausibility. For example, consider the instruction, “set up a dining table.” A vision-language model might initially generate poses where the chairs overlap with the table. Our differentiable spatial relations are designed to adjust these poses during optimization—preventing overlaps—while maintaining the essential semantics, such as “chairs should be positioned near the dining table in a dining setup.” To design a set of spatial relations that can capture a wide range of semantics for indoor scenes, we devise five expressive spatial relations: two positional objectives (i.e., distance, on_top_of), two orientational objectives (i.e., align_with, point_towards), and one wall-related objective (i.e., against_wall) that pertains to both the position and orientation of an asset. Note that our spatial relations do not rely on a fixed reference frame, unlike classic spatial relations like in_front_of and left_of. Each spatial relation corresponds to a differentiable objective function with optional parameters that VLMs can choose to impose on object poses; for example, a distance imposes a higher loss if objects are outside a specified distance range, and the VLMs can decide the lower and upper bound for this function. Table 1 presents the notations and meanings of these spatial relations. Formally, we denote a set of spatial relations derived from a scene layout representation as $\mathcal{R}$.
@@ 4.2 Generating Scene Layout Representation with Vision-Language Models
Self-Consistent Decoding. One key challenge is that VLMs struggle with spatial planning; while they may successfully generate spatial relations for pairs of objects, they tend to fail at accounting for overall layout coherence. We hypothesize that self-consistent spatial relations (i.e., ”the spatial relations that are also satisfied in the estimated numerical poses of the objects”) represent the most critical semantics to preserve during optimization when the object poses are adjusted for better physical plausibility. Thus, we introduce self-consistent decoding for our scene layout representation. Different from standard self-consistency [30], which selects the most consistent answer from multiple reasoning paths following the same format, we require the two distinct but mutually reinforcing representations predicted by the VLM to self-consistent. That is, we only retain the spatial relations satisfied with the predicted poses. After self-consistent decoding, we can formally describe the semantic part of the optimization loss as:
4.3 Differentiable Optimization
4.4 Finetuning VLMs with Scene Datasets
@@ 5.1 Experimental Setup
Evaluation Metrics. We evaluate generated 3D layouts on physical plausibility and semantic coherence with respect to the provided language instructions. Physical plausibility is measured using the Collision-Free Score (CF) and In-Boundary Score (IB). All assets are enforced to be placed, with remaining assets randomly placed if a method fails. Semantic coherence is assessed using Positional Coherency (Pos.) and Rotational Coherency (Rot.), measuring alignment with the input prompt. To evaluate semantic coherence across layouts without groundtruth, we use GPT-4o to score layouts based on top-down and side-view renderings and the language instructions, leveraging its effectiveness as a human-aligned evaluator in text-to-3D generation [35]. We also introduce the Physically-Grounded Semantic Alignment Score (PSA) to assess physical plausibility and semantic alignment, assigning 0 if assets cannot be feasibly placed. Scores range from 0 to 100, with higher scores indicating better performance.
@@ 5.2 Benchmark Performance
w/o Visual 67.2 $\pm$ 1.1 94.9 $\pm$ 0.8 79.1 $\pm$ 1.3 74.0 $\pm$ 1.2 48.2 $\pm$ 1.5
@@
Table 3: Ablation. Removing visual input impedes spatial planning, while omitting self-consistency and optimization negatively impacts both physical feasibility and layout coherence.
@@ 5.3 Ablation Study
We conduct an ablation study to assess the key components of our approach. In w/o Visual, the VLM was replaced with an LLM. The ablation w/o Self-Consistency removes the step of validating predicted constraints with raw poses, while w/o Const. placed objects based solely on predicted poses, without performing optimization. As shown in Table 3, the results confirm the effectiveness of our method’s design. Although w/o Visual still uses the same optimization process, which improves physical feasibility, we observe a clear advantage in using a VLM. This likely stems from the VLM’s ability to leverage rendered scenes and asset views to enhance spatial planning (e.g., arranging groups of objects in distinct regions). The comparison with w/o Const. confirms that VLMs alone were insufficient; our scene representation—consisting of both numerical object poses and spatial relations-is essential for generating practical layouts. Finally, we observe that self-consistency takes advantage of our scene representation to help ensure that placements met both physical and semantic requirements.
@@ 5.4 Finetuning LayoutVLM
In this experiment, we investigate whether fine-tuning pre-trained VLMs with our scene layout representation improves physical feasibility and semantic coherence. Specifically, we compare this approach with fine-tuning VLMs to directly predict numerical poses. Since the 3D-Front training data consist of typical household room types, we evaluated the fine-tuned models on test cases within the residential category (i.e., bedroom, living room, and dining room). As these test cases include unseen 3D assets from Objaverse and even new object categories, this experiment also assessed generalization capability.
@@ A.2 Differentiable Spatial constraint
The ”Against-Wall” objective consists of two components: (a) the sum of distances from the object’s corners to the wall, and (b) a term that encourages the object to point away from the wall. Let $c_{i}^{(k)}$ be the $k$-th corner of the $i$-th object, which is calculated based on the object’s $x$-$y$ position $(x_{i},y_{i})$, rotation, and bounding box size $b_{i}$. The loss function is:
@@ A.3 Self-Consistent Decoding
We propose self-consistent decoding to address the challenge of maintaining layout coherence in VLM-generated spatial plans. Our main hypothesis is that preserving self-consistent spatial relations—those that align with the estimated numerical poses of objects—is essential for ensuring semantic and physical plausibility during optimization. During implementation, we simplify the decoding process by enforcing that each asset maintains at most one orientational constraint, either to “point towards” or “align with” another asset. Additionally, the spatial relation ”on top of” is excluded from the self-consistency decoding, as we empirically observe that “on top of ” relations are almost accurately and reasonably predicted by our model; thus, enforcing self-consistency is unnecessary.
@@ B.1 Generating Test Cases
Requirement1:Bespecifc␣␣abouthowyoudescribetheobjects,whileusingassimpleenglishaspossible.Eachobjectshouldtrytobearoundtwowords,including

### v2

@@ Abstract
Spatial reasoning is a fundamental aspect of human cognition, enabling intuitive understanding and manipulation of objects in three-dimensional space. However, Large Language Models (LLMs) struggle with simple tasks such as arranging 3D assets in space according to open-ended language instructions, particularly in dense and physically constrained environments. We introduce LayoutVLM, a framework and scene layout representation that exploits the semantic knowledge of Vision-Language Models (VLMs) and supports differentiable optimization to ensure physical plausibility. LayoutVLM employs VLMs to generate two mutually reinforcing representations from visually marked images, and a self-consistent decoding process to improve VLMs spatial planning. Our experiments show that LayoutVLM addresses the limitations of existing LLM and constraint-based approaches, producing physically plausible 3D layouts better aligned with the semantic intent of input language instructions. We also demonstrate that fine-tuning VLMs with the proposed scene layout representation extracted from existing scene datasets can improve their reasoning performance.
@@ 4.1 Scene Layout Representation
Differentiable Spatial Relations. The goal of these spatial relations is twofold: (a) to capture the semantics of the input language instruction and (b) to preserve these semantics during optimization for physical plausibility. For example, consider the instruction, “set up a dining table." A vision-language model might initially generate poses where the chairs overlap with the table. Our differentiable spatial relations are designed to adjust these poses during optimization—preventing overlaps—while maintaining the essential semantics, such as “chairs should be positioned near the dining table in a dining setup." To design a set of spatial relations that can capture a wide range of semantics for indoor scenes, we devise five expressive spatial relations: two positional objectives (i.e., distance, on_top_of), two orientational objectives (i.e., align_with, point_towards), and one wall-related objective (i.e., against_wall) that pertains to both the position and orientation of an asset. Note that our spatial relations do not rely on a fixed reference frame, unlike classic spatial relations like in_front_of and left_of. Each spatial relation corresponds to a differentiable objective function with optional parameters that VLMs can choose to impose on object poses; for example, a distance imposes a higher loss if objects are outside a specified distance range, and the VLMs can decide the lower and upper bound for this function. Table 1 presents the notations and meanings of these spatial relations. Formally, we denote a set of spatial relations derived from a scene layout representation as $\mathcal{R}$.
@@ 4.2 Generating Scene Layout Representation with Vision-Language Models
Self-Consistent Decoding. One key challenge is that VLMs struggle with spatial planning; while they may successfully generate spatial relations for pairs of objects, they tend to fail at accounting for overall layout coherence. We hypothesize that self-consistent spatial relations (i.e., "the spatial relations that are also satisfied in the estimated numerical poses of the objects") represent the most critical semantics to preserve during optimization when the object poses are adjusted for better physical plausibility. Thus, we introduce self-consistent decoding for our scene layout representation. Different from standard self-consistency [30], which selects the most consistent answer from multiple reasoning paths following the same format, we require the two distinct but mutually reinforcing representations predicted by the VLM to self-consistent. That is, we only retain the spatial relations satisfied with the predicted poses. After self-consistent decoding, we can formally describe the semantic part of the optimization loss as:
4.3 Differentiable Optimization
4.4 Finetuning VLMs with Scene Datasets
@@ 5.1 Experimental Setup
Evaluation Metrics. We evaluate generated 3D layouts on physical plausibility and semantic coherence with respect to the provided language instructions. Physical plausibility is measured using the Collision-Free Score (CF) and In-Boundary Score (IB). All assets are enforced to be placed, with remaining assets randomly placed if a method fails. Semantic coherence is assessed using Positional Coherency (Pos.) and Rotational Coherency (Rot.), measuring alignment with the input prompt. To evaluate semantic coherence across layouts without groundtruth, we use GPT-4o to score layouts based on top-down and side-view renderings and the language instructions, leveraging its effectiveness as a human-aligned evaluator in text-to-3D generation [35]. We also introduce the Physically-Grounded Semantic Alignment Score (PSA) to assess physical plausibility and semantic alignment, assigning 0 if assets cannot be feasibly placed. PSA is calculated simply the GPT-4o rating weighted by physical plausibility. Scores range from 0 to 100, with higher scores indicating better performance.
@@ 5.2 Benchmark Performance
User GPT-4o Position Rotation PSA Position Rotation PSA LayoutGPT 1.91 1.86 2.77 2.00 1.82 2.83 Holodeck 3.44 3.50 3.10 3.20 3.43 3.10 I-Design 2.86 2.91 2.64 2.89 2.86 2.62 LayoutVLM 1.79 1.73 1.50 1.91 1.89 1.45
Table 3: Average ranks based on user ratings and GPT4-o scores.
Position Rotation PSA Within Users 0.51 0.57 0.50 Users With GP4-o 0.49 0.61 0.46
Table 4: User-User and User-GPT4o Alignment/Agreement
(a) LayoutVLM
(b) w/o self-consistency decoding
Figure 6: Comparison when w/ and w/o self-consistency decoding.
5.3 Human Eval to validate our metrics
We follow our baseline I-Design, evaluating layouts using physical plausibility metrics and GPT-4o-rated scores. To further evaluate the alignment between GPT-4o ratings and human ratings, we conducted a user study. Following prior work [35], we recruited five graduate students to rank the methods based on position, orientation, and overall performance. They were given the same instructions used as prompts for the GPT-4o evaluator. We collected 495 ratings for each method and metric pair. We converted GPT-4o scores to rankings and computed Kendall’s Tau [kendall1938new] in Table 4, showing strong user-user agreement and user-GPT agreement. Table 3 reports the average rankings based on user and GPT-4o ratings, demonstrating strong correlation with the results in our proposed metrics.
@@ 5.3 Human Eval to validate our metrics
Ablating the key components in LayoutVLM
w/o Visual Image 67.2 $\pm$ 1.1 94.9 $\pm$ 0.8 79.1 $\pm$ 1.3 74.0 $\pm$ 1.2 48.2 $\pm$ 1.5
@@
Ablating the two types of visual marks
w/o Visual Asset Mark 79.8 $\pm$ 0.6 87.9 $\pm$ 3.9 77.7 $\pm$ 0.6 72.2 $\pm$ 0.8 52.2 $\pm$ 5.0
w/o Visual Coordinate 77.2 $\pm$ 2.0 88.9 $\pm$ 1.7 74.3 $\pm$ 1.0 68.9 $\pm$ 1.3 46.0 $\pm$ 2.7
w/o Any Visual Mark 74.2 $\pm$ 6.4 84.9 $\pm$ 3.1 74.5 $\pm$ 0.8 69.2 $\pm$ 0.3 43.0 $\pm$ 2.2
Ablating the proposed scene layout representation
w/o Numerical Init. 74.7 $\pm$ 4.0 86.6 $\pm$ 2.6 69.5 $\pm$ 0.1 65.4 $\pm$ 1.4 41.0 $\pm$ 1.7
w/o (Spatial) Const. 75.8 $\pm$ 2.5 14.1 $\pm$ 3.0 70.4 $\pm$ 1.3 67.3 $\pm$ 0.9 6.7 $\pm$ 1.6
Table 5: Ablation. Removing visual input impedes spatial planning, while omitting self-consistency and optimization negatively impacts both physical feasibility and layout coherence. We also ablate the importance of both types of visual marks and the two key components in our proposed scene layout representation.
@@ 5.4 Ablation Study
We conduct an ablation study to assess the key components of our approach. In w/o Visual, the VLM was replaced with an LLM. The ablation w/o Self-Consistency removes the step of validating predicted constraints with raw poses, while w/o Const. placed objects based solely on predicted poses, without performing optimization. As shown in Table 5, the results confirm the effectiveness of our method’s design. Although w/o Visual still uses the same optimization process, which improves physical feasibility, we observe a clear advantage in using a VLM. This likely stems from the VLM’s ability to leverage rendered scenes and asset views to enhance spatial planning (e.g., arranging groups of objects in distinct regions). The comparison with w/o Const. confirms that VLMs alone were insufficient; our scene representation—consisting of both numerical object poses and spatial relations-is essential for generating practical layouts. Finally, we observe that self-consistency takes advantage of our scene representation to help ensure that placements met both physical and semantic requirements. We also analyze the effects of two key components in our proposed scene layout representation. The results indicate that spatial constraints are crucial for ensuring physical validity, particularly in preventing object overlap, while a good numerical initialization contributes to better semantic scores.
@@ 5.5 Finetuning LayoutVLM
In this experiment, we investigate whether fine-tuning pre-trained VLMs with our scene layout representation improves physical feasibility and semantic coherence. Specifically, we compare this approach with fine-tuning VLMs to directly predict numerical poses. Since the 3D-Front training data consist of typical household room types, we evaluated the fine-tuned models on test cases within the residential category (i.e., bedroom, living room, and dining room). As these test cases include unseen 3D assets from Objaverse and even new object categories, this experiment also assessed generalization capability. We extract the ground-truth layouts from 3D scenes by selecting defined spatial relations higher than predetermined thresholds. The language instructions are generated based on the type of rooms annotated in the 3D scene dataset. We finetune the open-sourced LLaVA-NeXT-Interleave [33] model via LoRA and GPT4-o via the publicly available finetuning API.
@@ A.2 Differentiable Spatial constraint
The "Against-Wall" objective consists of two components: (a) the sum of distances from the object’s corners to the wall, and (b) a term that encourages the object to point away from the wall. Let $c_{i}^{(k)}$ be the $k$-th corner of the $i$-th object, which is calculated based on the object’s $x$-$y$ position $(x_{i},y_{i})$, rotation, and bounding box size $b_{i}$. The loss function is:
@@
Optimization Details
We use Adam optimizer and Exponential LR scheduler with a decay factor $0.96$. Each optimization runs for 400 steps with projection back to the boundary every 100 iterations. Optimizing a scene with  40 assets takes 1–5 minutes on a single GPU, 5-10 GPT-4o calls (i.e., one per group), varying based on the VLM-defined optimization problem.
@@ A.3 Self-Consistent Decoding
We propose self-consistent decoding to address the challenge of maintaining layout coherence in VLM-generated spatial plans. Our main hypothesis is that preserving self-consistent spatial relations—those that align with the estimated numerical poses of objects—is essential for ensuring semantic and physical plausibility during optimization. During implementation, we simplify the decoding process by enforcing that each asset maintains at most one orientational constraint, either to “point towards" or “align with" another asset. Additionally, the spatial relation "on top of" is excluded from the self-consistency decoding, as we empirically observe that “on top of " relations are almost accurately and reasonably predicted by our model; thus, enforcing self-consistency is unnecessary.
@@ B.1 Generating Test Cases
Requirement 1: Be specifc about how you describe the objects, while using as simple english as possible. Each object should try to be around two words, including